# Semantic Segmentation , Self-Driving Car , Computer vision , BDD100K , Light-weight model , Embedded devices

### v4

@@ 4.2.2 Drivable Area Segmentation & Lane Detection Result
In this section, we proceed to compare our models with other models that undertake the same tasks of drivable area segmentation and lane detection. To ensure fairness, we will not compare with single-task models or models that perform an additional task such as object detection. All our experimental results in this section are presented in Fig 1 and Table 3. The provided table offers a comprehensive overview of the performance of models in drivable area segmentation and lane detection. The TwinLiteNet+ models, particularly the Nano, Small, and Medium configs, stand out for their lightweight architecture and high-speed performance. Despite TwinLiteNet achieving only about 87.3% mIoU, it has the lowest requirements with just 0.57G FLOPs and 0.03M parameters. The larger configs, like TwinLiteNet and TwinLiteNet, achieve a significant improve in mIoU and IoU metrics, with TwinLiteNet reaching 92.9% in mIoU and 34.2% in IoU, surpassing all current models and approaching top-tier performance while still maintaining optimal parameter and FLOPs usage. Other models like DeepLabV3+ [40], SegFormer[41], and YOLOP [10] also perform well but require more resources. This highlights the advantage of TwinLiteNet+ in terms of performance and resource efficiency. Notably, TwinLiteNet not only improves in mIoU but also accuracy and IoU for lane detection, achieving the highest levels compared to other models. Overall, the table demonstrates the trade-off between performance and resource utilization. While larger models like TwinLiteNet offer high performance, smaller ones like Nano and Small have the advantage of speed and low resource requirements, suitable for applications needing quick and efficient responses. These results provide critical information for researchers and engineers when selecting and optimizing models for practical applications.
We visually compare the TwinLiteNet+ with the TwinLiteNet model and evaluate the performance not only in daylight conditions but also at night. As we are concerned, intense sunlight or low-light environments can affect the driver’s visibility. Similarly, it impacts the model’s performance. This poses challenges for the performance of drivable area and lane segmentation. In Figure 6, the various configs of TwinLiteNet+ are compared to their predecessor, TwinLiteNet, in the two tasks of drivable area and lane segmentation. Overall, the TwinLiteNet and TwinLiteNet models perform significantly better than the basic TwinLiteNet, showing superior architecture and excelling in handling complex road constructions and diverse environmental conditions. As seen in Figure 6(b) and Figure 6(c), TwinLiteNet shows more misclassified regions and unclear boundaries compared to a more consistent and accurate segmentation of roads of the aforementioned. Conversely, our ultra-lightweight model config designed for situations with limited computational resources did not perform as expected in complex road structures, however, still reflects an effort in recognizing lanes and roads as in Figure 6(c) and Figure 6(d).
(a)Ground truth
(b)Drivable Area Segmentation
(c)Directly & Alternative Area segmentation
Figure 7: Examples of Ground truth visualization for Directly & Alternative Area segmentation task.
4.2.3 Directly Alternative Area segmentation results
Figure 8: Results visualization of TwinLiteNet for Directly Alternative Area segmentation. Red regions are directly drivable area, the blue ones are alternative and the lanes are green.
Table 5: Detailed analysis of TwinLiteNet in different environmental conditions.
Photometric scene
Number
Drivable Area
Lane
mIoU (%)
IoU (%)
Weatherconditions
Clear
5,346
93.1
33.8
Overcast
1,239
93.7
35.8
Undefined
1,157
93.2
35.4
Snowy
769
91.6
31.6
Rainy
738
90.2
32.0
Partly Cloudy
738
93.3
35.3
Foggy
13
92.4
29.6
Periodsof day
Daytime
5,258
93.2
35.1
Night
3,929
92.5
32.9
Dawn/Dusk
778
92.9
33.9
Undefined
35
86.7
29.3
Trafficscenes
City street
6,112
92.8
34.4
Highway
2,499
93.3
33.8
Residential
1,253
92.8
34.4
Undefined
53
90.3
31.0
Parking lot
49
88.1
25.1
Tunnel
27
89.8
28.9
Gas station
7
88.8
14.2
In this study, we expand our research to include a model that focuses on the division of directly drivable and alternative areas. This approach differs from the standard drivable area segmentation, which usually combines these two distinct areas into a single area. Figure 7 illustrates the main difference between the two segmentation tasks, depicted in Figure 7(b) and Figure 7(c), where Figure 7(c) is segmented into direct and alternative areas. This division of regions allows vehicles to differentiate between navigable regions and alternative routes, making it easier to change lanes or maneuver around obstacles in the region. When performing this segmentation, the controllable segmentation decoder block output has three channels instead of two. Therefore, the output of the model for the driving area segment is . We train the model using two designs: One with a single decoder for direct and alternative driving area segmentation and another with two decoders for multi-task models (Lane segmentation and Directly Alternative Area segmentation).
Table 6: Performance benchmarking of TwinLiteNet+ across different Quantization approaches: Floating Point 16-bit (FP16) and Integer 8-bit (INT8). INT8 is implemented with two methods: Post-Training Static Quantization (PTSQ) and Quantization-Aware Training (QAT).
Config
Drivable Area
Lane
mIoU (%)
Accuracy (%)
IoU (%)
FP32
FP16
INT8
FP32
FP16
INT8
FP32
FP16
INT8
PTSQ
QAT
PTSQ
QAT
PTSQ
QAT
Nano
87.3
87.3
0.0
86.0
1.3
86.9
0.4
70.2
70.1
0.1
67.2
3.0
69.9
0.3
23.3
23.2
0.1
21.8
1.5
23.2
0.1
Small
90.6
90.6
0.0
88.5
2.1
90.2
0.4
75.8
75.7
0.1
74.6
1.2
75.6
0.2
29.3
29.3
0.0
28.3
1.0
28.8
0.5
Medium
92.0
92.0
0.0
91.6
0.4
91.9
0.1
79.1
79.1
0.0
77.1
2.0
78.6
0.5
32.3
32.2
0.1
31.5
0.8
31.9
0.4
Large
92.9
92.9
0.0
92.7
0.2
92.7
0.2
81.9
81.9
0.0
80.2
1.7
81.4
0.5
34.2
34.2
0.0
33.6
0.6
34.0
0.2
Table 7: Latency and Power Consumption benchmarking of TwinLiteNet+ on Embedded Devices (Jetson Xavier and Jetson TX2) in different Quantization methods.
Config
Jetson Xavier
Jetson TX2
INT8
FP16
FP32
FP16
FP32
Latency(ms)
Nano
7.553±0.046
9.155±0.092
10.303±0.222
22.268±0.156
26.374±0.058
Small
10.286±0.093
12.861±0.941
15.662±0.055
33.818±0.134
41.961±0.035
Medium
15.548±0.083
20.460±0.067
27.985±0.166
61.785±0.015
84.631±0.096
Large
29.102±0.107
43.089±0.198
69.150±0.382
153.24±0.094
218.150±0.099
Power(Watt)
Nano
9.657±2.034
10.029±2.267
11.790±2.089
3.647±0.637
4.097±0.603
Small
10.926±1.924
11.086±2.385
12.732±2.505
4.043±0.463
4.513±0.376
Medium
12.427±2.086
13.585±1.832
15.496±1.549
4.496±0.202
4.747±0.122
Large
14.666±1.632
15.980±1.187
17.710±0.594
4.883±0.057
5.019±0.051
Table 4 provides a comprehensive and detailed overview of the performance of various segmentation models applied to direct region segmentation and replacement. In addition to the evaluation mIoU metric, in this study, we also assess the Pixel Accuracy (PA) and Mean Pixel Accuracy (mPA), facilitating comparison with previous research. In this experiment, we compare the TwinLiteNet model after modifications to demonstrate the responsiveness of our model to this task, and we name it as TwinLiteNet to differentiate it from the Large one. Additionally, we experiment with the TwinLiteNet model when performing multi-tasks and the TwinLiteNet with only one decoding block for the driving area segmentation. Notably, the IBN-NET model from 2020 demonstrates a significant improvement in mIoU, indicating its robustness in capturing intricate area details compared to its predecessors. The TwinLiteNet models, in both multitask and single-task configurations, stand out with high performance across all three metrics, with our TwinLiteNet model achieving the highest mPA of 92.1%, surpassing previous studies. This comprehensive assessment underscores the advancements in segmentation techniques and highlights the TwinLiteNet model’s potential to improve accuracy and detail capture in image segmentation tasks. We present some results of the TwinLiteNet model in Figure 8 to emphasize its proficiency in differentiating between directly drivable and alternative paths under various lighting conditions, from daylight to nighttime scenarios.
4.2.4 Panoptic driving perception results
The BDD100K dataset provides a vast collection of annotations for weather conditions, periods of day, and various scenes, ranging from urban streets and expansive highways to residential neighborhood areas. This dataset’s distribution mirrors the diverse content found in images, offering a rich landscape for exploration in image domain adaptation. This eclectic mix of images serves as a compelling avenue for research in the context of self-driving vehicles. Consequently, we embark on a thorough quantitative assessment of our TwinLiteNet+ configs, evaluating their adaptability and performance across distinct environmental contexts and scenarios. The results presented in Table 5 demonstrate that our model, TwinLiteNet+, exhibits proficient inference capabilities across diverse environments and conditions. These findings underscore the model’s adaptability and generalization ability in various conditions and scenarios related to autonomous driving applications.
4.3 Deployment
Table 8: TwinLiteNet+ model’s evaluation between Multitasking Learning versus Single-Task Learning approach.
Method
Drivable Area
Lane
Parameter
GFLOPs
mIoU (%)
Acc (%)
IoU (%)
Drivable (only)
92.6
1.91M
16.97
Lane (only)
81.9
34.4
1.91M
16.97
Multi-task
92.9
0.3
81.9
0.0
34.2
0.2
1.94M
0.03
17.58
0.61
To demonstrate the practical applicability of our TwinLiteNet+ model on hardware with limited computational capabilities, we conduct experiments using various inference data types, including Floating Point 32-bit (FP32), Floating Point 16-bit (FP16), and Integer 8-bit (INT8). For the INT8 data type, we employ quantization techniques, specifically Post-Training Static Quantization (PTSQ) and Quantization-Aware Training (QAT). With QAT, instead of training the model from scratch, we leverage the pre-trained technique and implemented QAT over 10 epochs. Table 6 demonstrates that TwinLiteNet+ maintains impressive performance using different quantization methods. Notably, the Quantization-Aware Training technique significantly mitigates the performance degradation commonly seen when transitioning from FP32 to INT8, which is crucial for applications with limited computational capabilities. The slight reduction in accuracy for tasks like drivable area and lane segmentation indicates a balance between accuracy and computational efficiency.
Table 9: TwinLiteNet+ Model’s evaluation between using ESP and DESP module. ✔ indicates that the model uses the DESP Module instead of ESP, and ✘ is the opposite.
DESP
Drivable Area
Lane
Param
FLOPs
mIoU (%)
IoU (%)
Nano
87.2
23.6
0.03M
0.57G
87.3
23.3
0.03M
0.57G
Small
90.7
29.2
0.15M
1.40G
90.6
29.3
0.12M
1.40G
Medium
92.3
32.6
0.62M
5.35G
92.0
32.3
0.48M
4.63G
Large
92.9
34.2
2.78M
22.19G
92.9
34.2
1.94M
17.58G
Furthermore, we implement our model on several embedded devices such as Jetson Xavier and Jetson TX2. TensorRT-FP32, FP16 are used respectively for inference and then measured the inference latency and power consumption on each device. With TensorRT-INT8, only Jetson Xavier is measured, as the Jetson TX2 isn’t supported yet. To minimize measurement discrepancies, we conduct five iterations, with each iteration performing 100 inferences. Results from Table 7 show that using the TwinLiteNet+ model on embedded devices like Jetson Xavier and Jetson TX2 yields promising results in terms of inference latency and power consumption. Specifically, the Nano and Small configs of the model exhibited low latency and reasonable power consumption on both devices with TensorRT-INT8, highlighting the model’s performance optimization capability in a limited computational environment. Although the Medium and Large configs have higher latency and power consumption, they still maintain an acceptable level, demonstrating the model’s flexibility and scalability. These results not only validate the impressive performance of TwinLiteNet+ but also underscore its practical potential in real-world applications, particularly in the field of autonomous vehicles and smart embedded systems.
Table 10: TwinLiteNet+ Model’s evaluation with different experimental settings.
Tversky Loss
Multiple Decoder
PCAA
EMA
Drivable Area
Lane
#Param.
GFLOPs
mIoU (%)
Acc (%)
IoU (%)
91.2
78.1
25.9
1.84M
16.83
91.3
0.1
80.9
2.8
29.3
3.4
1.84M
0.00
16.83
0.00
92.4
1.2
81.2
3.1
33.5
7.6
1.87M
0.03
17.42
0.59
92.7
1.5
81.7
3.6
33.9
8.0
1.94M
0.1
17.5
0.67
92.9
1.7
81.9
3.8
34.2
8.3
1.94M
0.1
17.5
0.67

### v5

@@ 4.2.2 Drivable Area Segmentation & Lane Detection Result
We evaluate the performance of our TwinLiteNet+ models versus state-of-the-art multitask models that jointly address drivable area segmentation and lane segmentation. For a fair comparison, we exclude single-task models and those incorporating additional tasks such as object detection. Table 3 summarizes the results along with computational complexity (FLOPs) and model size (number of parameters).
The original TwinLiteNet achieves 91.3% mIoU and 31.1% IoU with only 3.9G FLOPs and 0.44M parameters, offering a strong efficiency–accuracy trade-off among lightweight models. This establishes it as a solid baseline in terms of performance and model compactness. Compared to this baseline, the results demonstrate that TwinLiteNet+ consistently outperforms other lightweight multitask models across all metrics. In particular, the TwinLiteNet configuration achieves the highest scores in both tasks-92.9% mIoU for drivable area segmentation and 34.2% IoU for lane detection-outperforming strong baselines such as SegFormer, Sparse U-PDP, and IALaneNet. Importantly, TwinLiteNet accomplishes this with a significantly smaller model size (1.94M parameters) and moderate FLOPs (17.58G) compared to heavyweight counterparts like IALaneNet or DeepLabV3+, making it highly competitive for both accuracy and efficiency. While the TwinLiteNet model exhibits lower accuracy (mIoU of 87.3%, IoU of 23.3%), it offers ultra-efficient inference, requiring just 0.57G FLOPs and 0.03M parameters, making it suitable for highly resource-constrained embedded platforms. The Small and Medium configurations offer balanced trade-offs: TwinLiteNet achieves 92.0% mIoU and 32.3% IoU with only 0.48M parameters-surpassing most existing models like YOLOP and even the lightweight BILane (which has 1.4M parameters but slightly lower IoU at 31.3%).
Overall, these results underscore the scalability and practicality of TwinLiteNet+. The model family provides a spectrum of choices-from highly compact and fast variants to larger, high-accuracy models-empowering developers to choose configurations that best match their target deployment environments. The clear performance advantage of TwinLiteNet+ across different configurations confirms its suitability for both high-performance and real-time embedded applications in autonomous driving.
4.3 Driving perception results in different environments
The BDD100K dataset offers diverse annotations for weather conditions, times of the day, and traffic scenes, providing a comprehensive foundation for evaluating model robustness under real-world driving scenarios. To this end, we conduct a detailed analysis of TwinLiteNet across multiple environmental contexts. As shown in Table 4, TwinLiteNet demonstrates stable performance across a wide range of environmental conditions. Notably, while conditions such as nighttime, Dawn/Dusk, and rainy weather are generally considered challenging due to poor lighting or visual degradation, the model still maintains strong segmentation performance (e.g., 92.5% mIoU for nighttime, 90.2% for rainy). This is attributed, in part, to the relatively high number of training samples available for these conditions in the BDD100K dataset, which allows the model to learn more generalized representations. In contrast, the performance tends to drop more significantly in rare scenarios, such as parking lots, tunnels, or gas stations, where the dataset provides few samples ( 500 samples for the training set). In these cases, the segmentation performance declines noticeably, suggesting that data scarcity remains a limiting factor in model generalization under extreme conditions. At the same time, despite having only around 130 samples, foggy scenes still achieve competitive performance-likely because their visual characteristics are not significantly different from those of other weather conditions, allowing the model to transfer learned features effectively.
These findings highlight the importance of both data diversity and representation balance in training robust multi-task models for autonomous driving. The qualitative results under different environmental conditions are further analyzed in Sec. 4.4
| Photometric scene | Training/Validation | Drivable Area | Lane |
| mIoU (%) | IoU (%) | ||
| Weather conditions | |||
| Clear | 37,344/5,346 | 93.1 | 33.8 |
| Overcast | 8,770/1,239 | 93.7 | 35.8 |
| Undefined | 8,119/1,157 | 93.2 | 35.4 |
| Snowy | 5,549/769 | 91.6 | 31.6 |
| Rainy | 5,070/738 | 90.2 | 32.0 |
| Partly Cloudy | 4,881/738 | 93.3 | 35.3 |
| Foggy | 130/13 | 92.4 | 29.6 |
| Times of the day | |||
| Daytime | 36,728/5,258 | 93.2 | 35.1 |
| Night | 28,108/3,929 | 92.5 | 32.9 |
| Dawn/Dusk | 5,027/778 | 92.9 | 33.9 |
| Undefined | 137/35 | 86.7 | 29.3 |
| Traffic scenes | |||
| City street | 43,516/6,112 | 92.8 | 34.4 |
| Highway | 17,379/2,499 | 93.3 | 33.8 |
| Residential | 8,074/1,253 | 92.8 | 34.4 |
| Parking lot | 377/49 | 88.1 | 25.1 |
| Undefined | 361/53 | 90.3 | 31.0 |
| Tunnel | 129/27 | 89.8 | 28.9 |
| Gas station | 27/7 | 88.8 | 14.2 |
4.4 Visualization
To provide a comprehensive evaluation, we conduct qualitative comparisons of different TwinLiteNet+ configurations against the baseline TwinLiteNet under various environmental conditions. Specifically, we visualize segmentation outputs under (1) different times of the day-including daytime, dawn/dusk, and nighttime-as shown in Figure 6; (2) Figure 7 illustrates the results across different weather conditions, including snowy, rainy, overcast, and foggy. and (3) Figure 8 shows the results under different traffic scenes, including residential areas, tunnels, highways, and parking lots. As shown in the figures, TwinLiteNet consistently delivers the most accurate segmentation results, followed by TwinLiteNet, which achieves significant improvements over the baseline TwinLiteNet while maintaining similar computational cost. The enhanced performance of TwinLiteNet+ is attributed not only to the integration of the Partial Class Activation Attention (PCAA) module-which facilitates more detailed spatial feature extraction through localized class-specific activations derived from Partial CAM-but also to the strengthened decoder design. The improved decoder incorporates additional convolutional layers after upsampling, allowing for better reconstruction of fine-grained features and contributing to smoother and more precise boundary delineation in segmentation outputs.
Furthermore, the qualitative results reveal that the baseline TwinLiteNet, as well as the lightweight configurations TwinLiteNet and , encounter challenges in complex environments. These include segmentation noise in cluttered scenes (e.g., residential, parking lot), inaccurate region segmentation (e.g., dawn/dusk), and missed detections of the small lanes (e.g., foggy weather). In contrast, the larger configurations-TwinLiteNet and TwinLiteNet-demonstrate superior performance in these scenarios due to their greater representational capacity and deeper architecture.
While TwinLiteNet and TwinLiteNet demonstrate lower segmentation accuracy, their lightweight architectures provide notable advantages in terms of computational efficiency and inference speed, making them highly suitable for resource-constrained embedded platforms. In contrast, higher-capacity configurations like TwinLiteNet and TwinLiteNet consistently achieve more accurate and fine-grained segmentation results, particularly in the aforementioned challenging scenarios, due to their enhanced spatial representation capabilities and deeper model capacity. These observations highlight the flexibility and scalability of the TwinLiteNet+ configs, enabling practical deployment across a wide range of autonomous driving applications-from high-performance systems to energy-efficient real-time use cases.
4.5 Practical Deployment Considerations
indicates a performance drop caused by quantization, whereas signifies that the accuracy remains unchanged. Config Drivable Area Lane mIoU (%) Acc (%) IoU (%) FP32 FP16 INT8 FP32 FP16 INT8 FP32 FP16 INT8 PTSQ QAT PTSQ QAT PTSQ QAT Nano 87.3 87.3 0.0 86.0 1.3 86.9 0.4 70.2 70.1 0.1 67.2 3.0 69.9 0.3 23.3 23.2 0.1 21.8 1.5 23.2 0.1 Small 90.6 90.6 0.0 88.5 2.1 90.2 0.4 75.8 75.7 0.1 74.6 1.2 75.6 0.2 29.3 29.3 0.0 28.3 1.0 28.8 0.5 Medium 92.0 92.0 0.0 91.6 0.4 91.9 0.1 79.1 79.1 0.0 77.1 2.0 78.6 0.5 32.3 32.2 0.1 31.5 0.8 31.9 0.4 Large 92.9 92.9 0.0 92.7 0.2 92.7 0.2 81.9 81.9 0.0 80.2 1.7 81.4 0.5 34.2 34.2 0.0 33.6 0.6 34.0 0.2
| Config | Jetson Xavier | Jetson TX2 | ||||
| INT8 | FP16 | FP32 | FP16 | FP32 | ||
| Latency (ms) | ||||||
| Nano | 07.553±0.046 | 09.155±0.092 | 10.303±0.222 | 022.268±0.156 | 026.374±0.058 | |
| Small | 10.286±0.093 | 12.861±0.941 | 15.662±0.055 | 033.818±0.134 | 041.961±0.035 | |
| Medium | 15.548±0.083 | 20.460±0.067 | 27.985±0.166 | 061.785±0.015 | 084.631±0.096 | |
| Large | 29.102±0.107 | 43.089±0.198 | 69.150±0.382 | 153.240±0.094 | 218.150±0.099 | |
| Power (Watt) | ||||||
| Nano | 09.657±2.034 | 10.029±2.267 | 11.790±2.089 | 003.647±0.637 | 004.097±0.603 | |
| Small | 10.926±1.924 | 11.086±2.385 | 12.732±2.505 | 004.043±0.463 | 004.513±0.376 | |
| Medium | 12.427±2.086 | 13.585±1.832 | 15.496±1.549 | 004.496±0.202 | 004.747±0.122 | |
| Large | 14.666±1.632 | 15.980±1.187 | 17.710±0.594 | 004.883±0.057 | 005.019±0.051 | |
4.5.1 Quantization
To evaluate the practical deployability of the TwinLiteNet+ model on hardware with limited computational resources, we conducted experiments using three common inference data types: 32-bit floating point (FP32), 16-bit floating point (FP16), and 8-bit integer (INT8). Among these, FP16 can be applied directly without requiring retraining or calibration. In contrast, quantization to INT8 demands additional processing to preserve accuracy.
Specifically, for Post-Training Static Quantization (PTSQ), we followed a standard calibration procedure by sampling 500 random images from the training set to estimate activation ranges. Based on this, we applied 8-bit quantization to both weights and activations. While this approach is straightforward to integrate post-training, it resulted in reduced segmentation accuracy, particularly in lane detection, where pixel-level precision is crucial. To mitigate this problem, we employed Quantization-Aware Training (QAT), which simulates quantization effects during training. Using the pytorch_quantization framework, we inserted fake quantization modules into the model’s computational graph. Rather than training from scratch, we fine-tuned a pre-trained model for an additional 10 epochs, enabling it to adapt to quantization noise while maintaining training stability.
As shown in Table 5, TwinLiteNet+ maintains strong segmentation performance under all quantization settings. While INT8 quantization involves a more complex process than FP16, QAT allows the model to retain accuracy comparable to FP32 while significantly improving inference efficiency and reducing model size-critical factors for deployment on resource-constrained embedded devices.
4.5.2 Inference Performance on Embedded Devices
Furthermore, we implement our model on several embedded devices such as Jetson Xavier and Jetson TX2. TensorRT-FP32, FP16 are used respectively for inference and then measured the inference latency and power consumption on each device. With TensorRT-INT8, only Jetson Xavier is measured, as the Jetson TX2 isn’t supported yet. To enable INT8 quantization, we adopt the Post-Training Static Quantization (PTSQ) method, as our primary focus is to assess inference efficiency rather than optimize quantization accuracy. In this process, we perform calibration using 500 randomly selected samples from the training set to estimate the activation ranges, following the setup described in Section 4.5.1. All inference experiments are performed using TensorRT with default settings, without any manual tuning or hardware-specific optimization. This setup aims to evaluate how well the model performs on various embedded devices. To minimize measurement discrepancies, we conduct five iterations, with each iteration performing 100 inferences. Results from Table 6 show that using the TwinLiteNet+ model on embedded devices like Jetson Xavier and Jetson TX2 yields promising results in terms of inference latency and power consumption. Specifically, the Nano and Small configs of the model exhibited low latency and reasonable power consumption on both devices with TensorRT-INT8, highlighting the model’s performance optimization capability in a limited computational environment. Although the Medium and Large configs have higher latency and power consumption, they still maintain an acceptable level, demonstrating the model’s flexibility and scalability. These results not only validate the impressive performance of TwinLiteNet+ but also underscore its practical potential in real-world applications, particularly in the field of autonomous vehicles and smart embedded systems.
4.6 Directly Alternative Area segmentation results
In this section, we expand our research to include a model that focuses on the division of directly drivable and alternative areas. This approach differs from the standard drivable area segmentation, which usually combines these two distinct areas into a single area. Figure 9 illustrates the main difference between the two segmentation tasks, depicted in Figure 9(b) and Figure 9(c), where Figure 9(c) is segmented into direct and alternative areas. This division of regions allows vehicles to differentiate between navigable regions and alternative routes, making it easier to change lanes or maneuver around obstacles in the region. When performing this segmentation, the controllable segmentation decoder block output has three channels instead of two. Therefore, the output of the model for the driving area segment is . We train the model using two designs: One with a single decoder for direct and alternative driving area segmentation and another with two decoders for multi-task models (Lane segmentation and Directly Alternative Area segmentation).
| Model | Directly & Alternative Area Segmentation | ||
| mIoU (%) | PA(%) | mPA(%) | |
| DeepLabv3+ [15] | 84.73 | – | – |
| PSPNet [7] | 84.13 | – | – |
| Pizzati et al. [28] | 82.62 | – | – |
| ShuDA-RFBNet [29] | 82.67 | – | – |
| Resnet+ASPP+FPN [30] | 84.58 | 97.09 | 91.14 |
| Resnet+FPN [30] | 82.70 | 96.51 | 91.42 |
| Resnet+ASPP+top-down [30] | 82.80 | 96.60 | 90.68 |
| Resnet+top-down [30] | 82.44 | 96.24 | 88.39 |
| Lee [31] | 83.34 | – | – |
| Sun et al. [32] | 83.01 | – | – |
| IDS-MODEL [33] | 83.63 | – | – |
| TwinLiteNet (single-task) | 83.9 | 96.9 | 92.0 |
| TwinLiteNet (multi-task) | 84.3 | 97.0 | 92.1 |
Table 7 provides a comprehensive and detailed overview of the performance of various segmentation models applied to direct region segmentation and replacement. In addition to the evaluation mIoU metric, in this study, we also assess the Pixel Accuracy (PA) and Mean Pixel Accuracy (mPA), facilitating comparison with previous research. In this experiment, we compare the TwinLiteNet model after modifications to demonstrate the responsiveness of our model to this task, and we name it as TwinLiteNet to differentiate it from the Large one. Additionally, we experiment with the TwinLiteNet model when performing multi-tasks and the TwinLiteNet with only one decoding block for the driving area segmentation. The TwinLiteNet models, in both multitask and single-task configurations, stand out with high performance across all three metrics, with our TwinLiteNet model achieving the highest mPA of 92.1%, surpassing previous studies and competitive in other metrics. This comprehensive assessment underscores the advancements in segmentation techniques and highlights the TwinLiteNet model’s potential to improve accuracy and detail capture in image segmentation tasks. We present some results of the TwinLiteNet model in Figure 10 to emphasize its proficiency in differentiating between directly drivable and alternative paths under various lighting conditions, from daylight to nighttime scenarios.